OpenStudy (anonymous):

find the tangents of the following curves x^2 + 16y^2 = 32 from (-6, 7/2). the answers are x+4y -8 and 41x + 4y + 232

OpenStudy (anonymous):

(-6, 7/2)

OpenStudy (anonymous):

IF YOU are having a tangent then consider 1st point a,b and slope of it from (-6,7/2) will be equal to to slope of parabola at that point... since its a parabola then you will get two line..

OpenStudy (anonymous):

okay thanks!!!

OpenStudy (anonymous):

no prob.. (:

OpenStudy (anonymous):

i don't think this is a parabola

OpenStudy (anonymous):

\[ x^2 + 16y^2 = 32\] \[2x+32yy'=0\] \[y'=-\frac{x}{16y}\]